# InterQnet: A Heterogeneous Full-Stack Approach to Co-designing Scalable Quantum Networks

### v1

@@ 4.5 Summary of Achievements and Open Challenges
For Yb atoms, we have demonstrated high-fidelity atom-photon entanglement directly in the telecom band—a key building block for remote entanglement. The main limitation today is photon collection efficiency. In the near term we are improving rates with parallel multiplexing; in the longer term we will couple atoms to optical cavities to enhance collection. In Er-doped TiO2, we have achieved a $1000\times$ narrowing of homogeneous linewidths at sub-Kelvin temperatures. This is critical for isolating single ions and achieving long coherence. Next steps include fabricating high-Q cavities and verifying performance with single emitters for telecom-compatible quantum memory. For superconducting devices, we are nearing completion of two key components: SNAIL-based readout resonators and piezo-optomechanical transducers with phononic crystal engineering. Upcoming work will focus on tuning nonlinearity and boosting mechanical Q to improve conversion efficiency. In frequency conversion, we have recently completed classical characterization of the devices, and we are now investigating noise sources—especially Raman noise—to support quantum-compatible conversion. Altogether, these efforts represent diverse but converging progress toward scalable, hardware-aware entanglement distribution across heterogeneous platforms.
@@
We are creating models for the above-mentioned key quantum devices in SeQUeNCe and applying numerical optimization methods to search for critical device parameters to maximize the fidelity of simulated networks. We successfully modeled the quantum transducer in SeQUeNCe [59]; the other key devices are works in progress. Via simulations, we will not only better understand the behaviors of current experiments but also obtain insights for future device designs and improvements. We will conduct “what-if” studies that will help capability development in future devices. For example, what if the dark count in the detectors is reduced by a factor of 10, or atomic operations can run an order of magnitude faster? Furthermore, we will develop specialized numerical optimization methods to optimize operation conditions for high-fidelity quantum communications in future large-scale quantum networks.
5 Error Management
@@
Different from qubit-based quantum computing, optical quantum communication relies on bosonic modes to generate entangled states for long-distance transmission, without the need for universal gate sets. Thus, we require methods that can tolerate photon loss, imperfect entanglement generation, and heterogeneous device noise, while remaining practical for near-term testbeds and scalable to future large networks. InterQnet adopts a layered approach to error management: (1) lightweight error detection via Pauli Check Sandwiching (PCS) and postselection, (2) resource-aware strategies that leverage gauge freedom to simplify correction, (3) entanglement distillation protocols that convert multiple noisy Bell pairs into fewer high-fidelity ones, and (4) advanced quantum LDPC (qLDPC) and bosonic codes that enable scalable, fault-tolerant entanglement purification. Together, these techniques form a hierarchy that links the device-level realities of Section 4 to the architectural protocols of Section 6, and they map naturally onto the two project goals: enabling practical demonstrations in InterQnet-Achieve and guiding at-scale design choices in InterQnet-Scale.
5.1 Pauli Check Sandwiching
@@
In previous work we investigated quantum error detection based on Pauli check sandwiching (PCS) [60], which shares many properties with quantum codes. In PCS the unitary circuit $U$ to be error mitigated is required to respect the symmetry $P_{L}UP_{R}=U$, where $P_{L}$ and $P_{R}$ are Pauli strings of the Pauli group. The control of Pauli operators is on ancillary qubits, and error detection is performed by measuring these ancillary qubits and postselecting on all zero outcomes. PCS provides protection on the targeted qubits and generally incurs less resource overhead than standard quantum error correction and detection codes do. From the point of view of quantum codes, PCS without checks on the ancillas is generally a distance 1 code (i.e., cannot detect arbitrary single-qubit errors), but we can detect arbitrary single-qubit errors on the targeted qubits by using the generators of the Pauli group as checks [60]. Additionally, there is a direct correspondence between quantum codes and entanglement purification [61, 62]. Thus, the performance of PCS in quantum networks is an important area of investigation.
The PCS scheme [60] for quantum networks performs PCS in a distributed fashion. In the simplest quantum repeater network, the PCS scheme can be executed as follows. Two nodes, $A$ and $B$, want to establish entanglement via a repeater $C$. AA and $B$ prepare GHZ-like states on their ends and send over two qubits each (one computation register qubit and one ancilla) to $C$ via an optical fiber link. These form the left end of the PCS scheme. The repeater $C$ then applies the right end of the PCS scheme along with entanglement-heralding measurements and then does postselection by performing measurements on the ancillary qubits. This can be enhanced by adding additional checks and ancillas. The novel contributions of our current work are as follows [63]. Assuming depolarizing errors, we derived the analytical equations for the postselected fidelity and postselection rate as a function of the noise for PCS $X$ checks and PCS X&ZX\&Z checks. For PCS $X$ checks we have
@@
where $c$ is the postselection probability, $F^{\prime}$ is the postselected Bell state fidelity, and $F$ is the initial fidelity (i.e., fidelity of the Bell state without PCS). Next, we simulated our scheme assuming depolarization noise on the optical fiber channel and witnessed fidelity gains. The fidelity gains also appeared in simulations with noisy gates. PCS can be enhanced by adding $Z$ checks on the ancilla qubits. We can apply this process recursively by also checking the new ancillas with $Z$ checks. We discovered that this process generates a family of distance 2 codes. The structure of this family is described in Fig. 4(a), where the red gates can be applied recursively. Since the code distance is fixed, only the first recursion of ancilla checks provides significant improvements in fidelity. We also simulated this method with noisy gates and attained fidelity improvements as shown in Fig. 4(b).
5.2 Quantum Resource Correction
@@
In certain cases we may be interested only in correcting certain properties of a quantum state. Intuitively, this leads to less resource overhead than standard quantum error correction codes (QECCs), which aim to correct the logical state exactly. Thus, we investigated correcting the amount of a quantum resource in [64]. In quantum networks, entanglement is the most commonly used resource. We derived necessary and sufficient conditions for logical coherence correction. While the conditions are equivalent to the QECC conditions, recovery is simplified. Resource-preserving operations can result in gauge freedom in the code space. For instance, for the $l_{1}$ norm measure of coherence [65] and incoherent basis in the standard logical basis, logical Pauli operations are gauge transformations. In contrast, in standard QECC, unwanted logical Pauli operations are logical errors. Thus, the optimization in decoders for stabilizer codes such as the surface code is simpler for resource correction than in QECC. This is an important result since the decoding problem is in the class of NP-complete or #P-complete problems, depending on whether degenerate decoding is used [66]. Importantly, fast decoders can have the same accuracy as generally more accurate decoders when applied to resource correction. These results extend to logical entanglement and other resources that have resource-preserving operations that have a similar structure. We also investigated extending the logical incoherent basis over the physical space and identified channels that preserve the coherence of any state in the code space. Consequently, stabilizer measurements can be combined in many situations.
5.3 Entanglement Distillation and Error Detection/Correction
@@
We optimize the entanglement purification protocol to reduce resource overhead while enhancing overall performance. In principle, entanglement distillation can asymptotically extract multiple copies of perfect Bell pairs, limited only by the distillable entanglement [67, 68]. In practice, with finite noisy entangled inputs, most purification protocols reduce noise and concentrate entanglement in the output but do not yield perfect Bell pairs. For instance, certain protocols require four noisy states as input to suppress depolarizing errors in the output state effectively. From an engineering perspective, these improved entangled states from finite-size protocols are valuable for constructing nontrivial applications.
@@
By exploiting the correspondence between quantum error correction/detection and entanglement distillation [69], we can devise a wide array of efficient entanglement purification protocols, assisted by one-way or two-way classical communication, based on quantum error-correcting codes. For example, Fig. 5(a) depicts a quantum circuit based on the $[[4,2,2]]$ code, also known to be equivalent to the Leung–Shor protocol [70]. Upon success, it transforms four noisy inputs into two purified entangled states with higher fidelity.
@@
To ensure practicality, we analyze the circuit implementations of entanglement purification on popular platforms. Real physical systems often impose additional constraints that hinder the efficient execution of the desired protocols. In Rydberg atom array systems, leveraging their unique features, we demonstrate the robustness and universality of efficient entanglement purification protocol implementations [71].
@@
5.4 qLDPC-Based Entanglement Distillation
@@
5.5 Summary of Achievements and Open Challenges
We have investigated several complementary approaches to error management. Pauli Check Sandwiching (PCS) improves fidelity through Pauli operator checks and postselection, though implementing the necessary entangling gates remains challenging. Quantum Resource Correction (QRC) exploits gauge freedom to construct resource-preserving recovery maps that simplify decoding relative to conventional QECC. Entanglement distillation protocols, designed for dual-species atom arrays and compiled into circuits with only global control, enable purification on hardware with limited addressability, though coupling and measurement performance must improve. At a more advanced level, quantum LDPC codes leverage sparse stabilizers to achieve fault-tolerant purification with high thresholds, while multimode bosonic codes such as GKP, combined with Petz recovery, can approach channel capacity [83, 84]; concatenating these methods provides a scalable path toward efficient entanglement distribution.
@@ 5.5 Summary of Achievements and Open Challenges
Together, these approaches provide a layered strategy. Lightweight PCS, QRC, and small-code distillation enable practical demonstrations in InterQnet-Achieve, while advanced qLDPC and bosonic codes inform scalability analyses in InterQnet-Scale.
6 Network Architecture and Protocols
@@
Fig. 6 illustrates a scalable quantum network organized using a three-plane abstraction. The infrastructure plane at the bottom represents all physical devices and fiber connections. The logical plane in the middle captures the repeater-chain abstraction once a network orchestrator has identified a suitable path between end nodes (e.g., A and B in dark green). The control and service plane at the top encompasses orchestration, routing, and user-facing services accessed through standard interfaces. We adopt this layered abstraction to structure our work on network architecture and protocols, enabling systematic study of scalability across distance, node and user counts, heterogeneity, and entanglement distribution rates, and providing a framework for integrating device models from Section 4 and error management methods from Section 5 into protocol-level evaluations.
6.1 Faithfully Simulating the Infrastructure Plane of Quantum Networks
Given the heterogeneous nature of InterQnet, SeQUeNCe has been designed to work with multiple architectural paradigms of the infrastructure plane. For this iteration of the work, we report simulation modes to tackle two important aspects of quantum communications: optical and atomic communications.
6.1.1 Simulating Optical Quantum Communications
The primary challenge in simulating fiber-optic quantum communication lies in the large Hilbert space required to model the many optical modes involved. Simulating open quantum systems with density matrices for such large systems is computationally expensive. Even with their sparsity, the exponentially increasing memory demands of sparse density matrices quickly make simulations intractable for systems with high local Hilbert space dimensions and multiple modes. To address this situation, we employed tensor networks to efficiently represent large state vectors as matrix product states [85].
Fig. 7 compares the memory requirements for simulating a polarization-based entanglement swapping experiment [86] using tensor networks versus sparse density matrices, showing a substantial reduction in memory usage as the Hilbert space dimension increases.

### v2

@@ IV-E Summary of Achievements and Open Challenges
Table I shows a summary of the current development status of InterQnet devices. For Yb atoms, we have demonstrated high-fidelity atom-photon entanglement directly in the telecom band—a key building block for remote entanglement. The main limitation today is photon collection efficiency. In the near term we are improving rates with parallel multiplexing; in the longer term we will couple atoms to optical cavities to enhance collection. This will bring us to the device performance parameter required for InterQnet-Achieve, nameky, 1 Hz heralded entanglement rate with fidelity >0.99. In Er-doped TiO2 we have achieved a $1000\times$ narrowing of homogeneous linewidths at sub-Kelvin temperatures. This is critical for isolating single ions and achieving long coherence. Next steps include fabricating high-Q cavities and verifying performance with single emitters for telecom-compatible quantum memory to achieve heralded entanglement rates of 1 Hz. For superconducting devices, we are nearing completion of two key components: frequency-tunable qubit readout resonators and high-efficiency, low-noise piezo-optomechanical transducers. Upcoming work will focus on engineering the nonlinearity and tunability of the readout resonator and boosting mechanical Q to improve the photon conversion efficiency of the transducer. These efforts will get us closer to the required >50% conversion efficiency with low added noise ($\lesssim 0.1$) for quantum operations for the InterQnet-Achieve demonstration. In frequency conversion, we have recently completed classical characterization of the devices, and we are now investigating noise sources—especially Raman noise—to support quantum-compatible conversion (i.e., efficiency >50% with negligible dark count noise). Altogether, these efforts represent diverse but converging progress toward scalable, hardware-aware entanglement distribution across heterogeneous platforms.
@@
To assist in the design and evaluation process, we are creating models for the above-mentioned key quantum devices in SeQUeNCe and applying numerical optimization methods to search for critical device parameters to maximize the fidelity of simulated networks. We have successfully modeled the quantum transducer, Yb atom node, and QFC in SeQUeNCe [29, 82]; the Er3+ device is work in progress. Via simulations, we will not only better understand the behaviors of current experiments but also obtain insights for future device designs and improvements. We will conduct “what-if” studies that will help capability development in future devices. For example, what if the dark count in the detectors is reduced by a factor of 10, or atomic operations can run an order of magnitude faster? Furthermore, we will develop specialized numerical optimization methods to optimize operation conditions for high-fidelity quantum communications in future large-scale quantum networks.
V Error Management
@@
Different from qubit-based quantum computing, optical quantum communication relies on bosonic modes to generate entangled states for long-distance transmission, without the need for universal gate sets. As a result, the quantum circuits involved in error management methods that can tolerate photon loss, imperfect entanglement generation, and heterogeneous device noise are usually simple and practical for near-term testbeds and scalable to future large networks. To aid in addressing challenges C2 and C3, InterQnet adopts a layered approach to error management: (1) lightweight error detection via Pauli check sandwiching (PCS) and postselection, (2) resource-aware strategies that leverage gauge freedom to simplify correction, (3) entanglement distillation protocols that convert multiple noisy Bell pairs into fewer high-fidelity ones, and (4) advanced quantum LDPC (qLDPC) and bosonic codes that enable scalable, fault-tolerant entanglement purification. Hardware-level realizations of these approaches have been partially demonstrated or proposed over the past decades. An early work [105] demonstrated primitive entanglement purification via postselection. Although quantum error-correcting codes (QECCs) have not been widely used in entanglement distillation, promising schemes that apply more scalable and powerful quantum codes have recently been proposed for superconducting qubits [45, 44] as well as for trapped ions and atom arrays [100, 13, 129, 16]. For bosonic quantum systems, error-correction protocols have been experimentally demonstrated in [90, 87, 113, 33, 101]. Together, these techniques form a hierarchy that links the device-level realities of Section IV to the architectural protocols of Section VI, and they map naturally onto the two project goals: enabling practical demonstrations in InterQnet-Achieve and guiding at-scale design choices in InterQnet-Scale.
V-A Pauli Check Sandwiching
@@
In previous work we investigated quantum error detection based on Pauli check sandwiching (PCS) [43], which shares many properties with quantum codes. PCS uses left and right Pauli checks to detect errors. It provides protection on the targeted qubits and generally incurs less resource overhead than standard quantum error correction and detection codes do. From the point of view of quantum codes, PCS is generally a distance 1 code (i.e., it cannot detect arbitrary single-qubit errors). Additionally, there is a direct correspondence between quantum codes and entanglement purification, since any qubit stabilizer codes can be converted to a one-way entanglement purification protocol due to the simulability of teleportation, and vice versa [4, 38]. Thus, the performance of PCS in quantum repeater networks is an important area of investigation for InterQnet-Scale.
The PCS scheme [43] for quantum networks performs PCS in a distributed fashion, where the left checks are performed by the origin node and the right checks are performed by the receiving node (see Alg. 1 for details). The left and right checks are chosen from the +1+1 elements of the Pauli group and must multiply to identity.
Algorithm 1 Network PCS for distributing a Bell pair
1:Nodes $A,B$; Pauli checks $\{P_{j}\}_{j=1}^{m}$; network channel $\mathcal{N}_{A\to B}$
2:Accept a (postselected) shared Bell pair on $(A,B)$ or Reject
3:AA prepares $|\Phi^{+}\rangle_{A\,F}$ $\triangleright$ $F$ is the half to be sent to $B$
[…]
5:  $a_{j}\leftarrow|+\rangle$ $\triangleright$ check ancilla prepared at AA
6:  apply controlled-$P_{j}$ with control $a_{j}$ and target $F$
7:end for
8:send $(F,\{a_{j}\}_{j=1}^{m})$ through the network: $(F,\{a_{j}\})\xrightarrow{\;\mathcal{N}_{A\to B}\;}(B,\{a_{j}\})$
[…]
10:  apply controlled-$P_{j}$ with control $a_{j}$ and target $B$
11:  measure $a_{j}$ in $X$ to obtain $s_{j}\in\{0,1\}$
12:end for
13:if $\forall j:\ s_{j}=0$ then
14:  return Accept
15:else
16:  return Reject
17:end if
The novel contributions of our current work [42] are as follows. Assuming depolarizing errors, we derived the analytical equations for the postselected fidelity and postselection rate as a function of the noise for PCS $X$ checks and PCS X&ZX\&Z checks. For PCS $X$ checks we have
@@
where $c$ is the postselection probability, $F^{\prime}$ is the postselected Bell state fidelity, and $F$ is the initial fidelity (i.e., fidelity of the Bell state without PCS). Next, we simulated our scheme assuming depolarization noise on the optical fiber channel and witnessed fidelity gains. The fidelity gains also appeared in simulations with noisy gates. PCS can be enhanced by adding $Z$ checks on the ancilla qubits. We can apply this process recursively by also checking the new ancillas with $Z$ checks. We discovered that this process generates a family of distance 2 codes. The structure of this family is described in Fig. 4(a), where the red gates can be applied recursively. Since the code distance is fixed, only the first recursion of ancilla checks provides significant improvements in fidelity. We also simulated this method with noisy gates and attained fidelity improvements as shown in Fig. 4(b). We vary along the x-axis the depolarization error rate of the flying qubits. The baseline without PCS is in pink. The blue, orange, and green lines are no recursion, one recursion, and two recursions, respectively. The dashed lines represent postselection rates. Since the code distance is fixed, the green and orange lines yield similar fidelities.
V-B Quantum Resource Correction
In certain cases we may be interested only in correcting certain properties of a quantum state. Intuitively, this leads to less resource overhead than standard QECCs, which aim to correct the logical state exactly. In quantum resource theory [22], restricted control defines a set of free operations $\mathcal{O}$ and free states $\mathcal{F}$. States outside $\mathcal{F}$ possess a quantum resource. The amount of resource in a state quantifies the degree to which it deviates from the free set, and thus how useful it is for tasks that cannot be accomplished by using only free states and operations. By definition, free states have zero resource, and the amount of resource cannot increase under free operations. This general notion underlies familiar quantities in quantum information such as entanglement, coherence, asymmetry, and nonlocality. Thus, in [17] we investigated error-correction schemes that restore a specified quantum resource while allowing other properties of the logical state to change (i.e., correcting the amount of a quantum resource).
@@ V-B Quantum Resource Correction
In quantum networks, entanglement is the most commonly used resource. We derived necessary and sufficient conditions for correcting the amount of a logical resource in a logical state. The resources considered are those that are left invariant by a group of unitary operations (e.g., local unitaries for entanglement). While the conditions are equivalent to the QECC conditions, recovery is simplified. Resource-preserving operations can result in gauge freedom in the code space. For instance, for the $l_{1}$ norm measure of coherence [7] and incoherent basis in the standard logical basis, logical Pauli operations are gauge transformations. In contrast, in standard QECC, unwanted logical Pauli operations are logical errors. Thus, the optimization in decoders for stabilizer codes (such as the surface code) is simpler for resource correction than in QECC. This is an important result since the general decoding problem is in the class of NP-complete or #P-complete problems, depending on whether degenerate decoding is used [35]. Importantly, fast decoders can have the same accuracy as generally more accurate decoders when applied to resource correction. We also investigated extending the logical incoherent basis over the physical space and identified channels that preserve the coherence of any state in the code space. Consequently, stabilizer measurements can be combined in many situations.
V-C Entanglement Distillation and Error Detection/Correction
@@
In principle, entanglement distillation can asymptotically extract multiple copies of perfect Bell pairs, limited only by the distillable entanglement [11, 1]. In practice, with finite noisy entangled inputs, the majority of purification protocols reduce noise and concentrate entanglement in the output but do not yield perfect Bell pairs. For instance, certain protocols require four noisy states as input to suppress depolarizing errors in the output state effectively. From an engineering perspective, these improved entangled states from finite-size protocols are valuable for constructing nontrivial applications.
@@
By exploiting the correspondence between quantum error correction/detection and entanglement distillation [3, 4, 38], we can devise a wide array of efficient entanglement purification protocols, assisted by one-way or two-way classical communication, based on quantum error-correcting codes. For example, Fig. 5(a) depicts a quantum circuit based on the $[[4,2,2]]$ code, also known to be equivalent to the Leung–Shor protocol [67]. Upon success, it transforms four noisy inputs into two purified entangled states with higher fidelity.
@@
To bridge the gap between theory and practice, we analyzed the circuit implementations of entanglement purification on popular platforms (e.g., neutral atom systems) in [69]. Real physical systems often impose additional constraints that hinder the efficient execution of the desired protocols. In Rydberg atom array systems, leveraging their unique features, we demonstrated the robustness and universality of efficient entanglement purification protocol implementations.
@@
V-D qLDPC-Based Entanglement Distillation
@@
V-E Summary of Achievements and Open Challenges
We have investigated several complementary approaches to error management. Pauli check sandwiching improves fidelity through Pauli operator checks and postselection, although implementing the necessary entangling gates remains challenging. Quantum resource correction (QRC) exploits gauge freedom to construct resource-preserving recovery maps that simplify decoding relative to conventional QECC. Entanglement distillation protocols, designed for dual-species atom arrays and compiled into circuits with only global control, enable purification on hardware with limited addressability, although coupling and measurement performance must improve. At a more advanced level, quantum LDPC codes leverage sparse stabilizers to achieve fault-tolerant purification with high thresholds, while multimode bosonic codes such as GKP, combined with Petz recovery, can approach channel capacity [133, 70]; concatenating these methods provides a scalable path toward efficient entanglement distribution.
@@ V-E Summary of Achievements and Open Challenges
Together, these approaches provide a layered strategy. Lightweight PCS, QRC, and small-code distillation enable practical demonstrations in InterQnet-Achieve, while advanced qLDPC and bosonic codes inform scalability analyses in InterQnet-Scale. To elaborate on the technical feasibility of employing these methods in ARQNET, we focus our discussion on atom and superconducting qubit platforms (see also Table I). Multiqubit control has been achieved by all these platforms with ¿99% fidelity, sufficient for QEC demonstrations [45, 13]. Quantum memory lifetimes (compared with typical classical communication time, millisecond or longer) impose an interesting trade-off. For atom platforms, the memory lifetime can be very long (approaching/exceeding a second), while the memory lifetime for superconducting qubits is often limited (approaching milliseconds for transmon qubits). Comparing these memory lifetimes with entanglement generation rates, atom platforms can achieve generation rates $\sim$1 Hz [71], close to the memory decoherence rate. For superconducting platforms, we can perform pitch-and-catch microwave excitations to generate entanglement between different dilution refrigerators (over 30 meters [117]). Hence, with good transmon qubits, we should be able to demonstrate entanglement distillation. For long-range entanglement generation, we will need M-O transduction before we can generate long-range entanglement of superconducting qubits. This is an active research area, as discussed in Section IV-C.
VI Network Architecture and Protocols
@@
This thrust directly tackles challenge C3 by conducting systems studies focused on architectural trade-offs for InterQnet-Scale. Fig. 6 illustrates a scalable quantum network organized using a three-plane abstraction. The infrastructure plane at the bottom represents all physical devices and fiber connections. The logical plane in the middle captures the repeater-chain abstraction once a network orchestrator has identified a suitable path between end nodes (e.g., A and B in dark green). The control and service plane at the top encompasses orchestration, routing, and user-facing services accessed through standard interfaces. We adopt this layered abstraction to structure our work on network architecture and protocols, enabling systematic study of scalability across distance, node and user counts, heterogeneity, while providing a framework for integrating device models from Section IV and error management methods from Section V into protocol-level evaluations.
VI-A Faithfully Simulating the Infrastructure Plane of Quantum Networks
Given the heterogeneous nature of InterQnet, SeQUeNCe has been designed to work with multiple architectural paradigms of the infrastructure plane. For this iteration of the work, we report simulation modes to tackle two important aspects of quantum communications: optical and atomic communications. The following subsections present updates on the development of software extensions to SeQUeNCe that will enable future large-scale studies.
VI-A1 Simulating Optical Quantum Communications
The objective of simulating optical quantum communications is to analyze the effect of noise on realistic communication experiments. However, modeling realistic systems requires simulating large Hilbert spaces. Density matrices soon become computationally intractable to model large systems. To address this issue, we use tensor networks [92] to simulate the sparsely entangled quantum states in quantum communications. Moreover, to model noise using tensor networks, we implement trajectory-based simulations to simulate open system dynamics.
Figure 7 compares the memory requirements for simulating a polarization-based entanglement swapping experiment [109] using tensor networks versus sparse density matrices, showing a substantial reduction in memory usage as the Hilbert space dimension increases.